Fermi National Laboratory
 

Operation Summary for January 21, 2011 - January 24, 2011

 

The Integrated Luminosity for the period from 1/17/11 to 1/24/11 was 70.18 pb-1
(inverse picobarns). NuMI reported receiving 8.61E18 protons on target during this same period.

Friday Jan. 21

The day shift began with Operations monitoring the Tevatron (TEV) store 8436 (247.63E30) and the Recycler stash (70.4E10), with the Antiproton Source (Pbar) stacking (15.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 (CALICE) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations monitored the accelerator complex.

During the evening shift, Operations disabled beam to T-978 at 6 PM. Operations terminated store 8436 at 9:48 PM, put the TEV into HEP shot setup, and established store 8437 at 10:52 PM with an average initial luminosity (AIL) of 354.35E30.

Saturday Jan. 22

The midnight shift began with Operations monitoring store 8437 (257.9E30) and the stash (69.1E10) and the Pbar stack (10.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 4 AM, Operations reported a LCW leak on Pbar Accumulator quadrupole magnet (A6Q10) that required a Pbar access at 7:32 AM to repair.

During the day shift, Water Group personnel reported at 10:44 AM that they had repaired magnet LCW leaks on A6Q10 and A6Q9.

During the evening shift, Operations terminated store 8347 at 4:41 PM, but shot setup was held off by noisy readbacks (T:IPROG and T:IRINGA). Operations put the TEV into shot setup at 7:45 PM and established store 8439 at 8:53 PM with an AIL of 353.05E30.

Sunday Jan. 23

The midnight shift began with Operations monitoring store 8439 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations monitored the accelerator complex.

During the day shift, Operations terminated store 8439 at 9:06 AM and put the TEV into shot setup, but the TEV quenched during the scrape at 10:15 AM. An interlocked air pressure readback on an expansion tank tripped the NuMI beam permit at 10:58 AM and at 12:22 PM; it reset.

During the evening shift, Operations put the TEV into shot setup at 5:24 PM and established store 8443 at 6:50 PM with an AIL of 225.38E30.

Monday Jan. 24

The midnight shift began with Operations monitoring store 8443 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations monitored the accelerator complex.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Linac personnel to switch from the I- Source to the H- Source today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/24/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory